Please help me with this problem:
a pharmacist wishes to mix a solution that is 6% minoxidil. She has on hand 100ml of a 1% solution and wishes to add some 8% solution to obtain the desired 6% solution. How much 8% solution should she add?

Answer by stanbon(75887)

a pharmacist wishes to mix a solution that is 6% minoxidil. She has on hand 100ml of a 1% solution and wishes to add some 8% solution to obtain the desired 6% solution. How much 8% solution should she add?
-----------------------------
Equation:
mino + mino = mino
0.01(100) + 0.08x = 0.06(100+x)
Multiply thru by 100 to get:
1*100 + 8x = 6*100 + 6x
2x = 500
x = 250 ml (amt. of 8% solution needed for the mixture)
==========================================================
Cheers,
Stan H.